Determine the magnitude of an electric field at r=0.01 m from the center of non-conducting sphere with a radius of R=0.001m possessing a charge of 7C which evenly distributed throughout the sphere. (r>R).


1
Expert's answer
2022-02-28T10:06:14-0500

Answer

Electric field is given by

"E=\frac{kQ}{r^2}\\=\frac{9*10^9*7}{0.01^2}\\=63*10^{13}N/m"